OpenStudy (secret-ninja):

Am I right? Medal. The product of (a + b)(a - b) is a perfect square trinomial. A. Sometimes B. Always C. Never I think its A.

OpenStudy (anonymous):

expand it out

OpenStudy (anonymous):

(a+b)(a-b) = a^2 - ba + ba - b^2 simplify it to get a^2 - b^2

OpenStudy (anonymous):

"Trinomial" is a term for a 3 term equation (a+b)(a-b)= a^2-ab+ab-b^2 combine like terms is there 3 terms?

OpenStudy (secret-ninja):

No, there is 4, so its a polynomial... So its C?

OpenStudy (anonymous):

Combining like terms there will be 2 terms so it would be a Binomial yes C

OpenStudy (anonymous):

i think this is called the difference of squares rules. It is worth remembering and no it is not a trinomial ever.

OpenStudy (secret-ninja):

Thanks ^-^

OpenStudy (cwrw238):

yep difference of 2 squares

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

you'll need it a lot when rationalizing denominators.